A sheet of metal is 15m long and 4m wide. There is a circle cut out of it.The diameter of the circle is 2m.How much metal is left?show work!
Answer by mananth(16949)

area of metal sheet = L*W
=15*4 = 60 sq. m
..
Area of circle = pi*r^2
r=2m
area = pi*2^2
12.56 sq. m
Area of plate - area of circle = remaining plate
= 60-32.56
=27.44 sq,m